# Brain-inspired Large Language Model

### v6

@@ Abstract
We introduce BriLLM, a brain-inspired large language model that redefines the foundations of generative language modeling. Departing from Transformer architectures, GPT frameworks, and traditional input-output constrained paradigms, BriLLM is built on the Signal Fully-connected flowing (SiFu) mechanism—a directed graph-based neural network design that enables full interpretability across all nodes, in contrast to conventional models limited to input-output interpretability. In this framework, tokens are represented as graph nodes, with signal flows—either randomly initialized or user-defined—propagating along paths following a "least resistance" principle. The next token to be generated emerges as the target of this signal flow. Theoretically, BriLLM supports infinitely long $n$-gram modeling, with model size decoupled from input and prediction length. Its signal propagation dynamics mimic human-like cognitive patterns, enabling recall activation and inherent multi-modal compatibility. We release initial Chinese and English BriLLM versions (4000 tokens, 32-dimensional nodes, 32-token sequence prediction capacity) with sizes  2B and  1B parameters, respectively, achieving performance comparable to GPT-1.  [^1] Code at: https://github.com/brillm05/BriLLM0.5  [^2] Models at: https://huggingface.co/BriLLM/BriLLM0.5
@@ 1 Introduction
Large language models (LLMs) have ignited unprecedented progress toward artificial general intelligence (AGI). Yet, state-of-the-art LLMs—rooted in Transformer architectures and GPT (Generative Pre-trained Transformer) training schemes—face fundamental limitations that may preclude them from achieving true AGI. Their scalability is constrained by computational bottlenecks, and their interpretability remains superficial, while their operational principles bear little resemblance to the human brain: a naturally evolved intelligence system refined over millions of years that offers a blueprint for AGI.
The Transformer (Vaswani et al., 2017) has become the backbone of modern LLMs, delivering remarkable generalization across tasks and scaling to increasingly sophisticated capabilities. However, the attention mechanism’s quadratic computational complexity with respect to sequence length creates intractable inefficiencies for long inputs. While efforts to develop linear attention variants or RNN-like Transformers aim to balance performance and cost, they merely mitigate rather than resolve this core bottleneck, as they retain attention-based architectures.
Compounding these challenges, Transformers suffer from limited parameter-level interpretability due to opaque self-attention dynamics and complex parameter interactions, functioning as "black boxes." Moreover, when deploying LLMs, GPT requires a quadratic increase in model size as input context length grows linearly. What is more, multi-modal large models based on Transformers or GPT depend heavily on multi-modal aligned data, making it far harder to build embodied intelligent systems that resemble human-like ones.
To address these critical limitations—which are not unique to Transformers or GPT but inherent to traditional machine learning—we rethink machine learning from its fundamental principles, integrating brain-inspired computational ideas. This is a straightforward line of reasoning: the brain, the seat of human intelligence, must provide key insights for overcoming the limitations of current machine learning frameworks as we move toward AGI. This rethinking has led to a novel machine learning mechanism termed Signal Fully-connected Flowing (SiFu) in this paper. SiFu learning is defined over a fully connected directed graph where all nodes are interpretable. Signaling in SiFu replaces the core prediction operations in traditional machine learning.
Building on the SiFu mechanism, we present BriLLM: a language model inspired by neural signal propagation in the brain. Unlike attention-based architectures, BriLLM functions on a bidirectional graph with nodes and edges, where each node (a neuron layer) represents a token, and edges—modeled as fully connected networks—encode token relationships. Signal propagation adheres to energy dynamics: tensors flow through the graph, with the next token determined by the path that maximizes signal energy.
Table 1 compares the two frameworks of existing machine learning and SiFu learning.
Table 1: Machine learning vs SiFu learning frameworks
GPT-LLM BriLLM
$\uparrow$ various "large" models various brain-like "large" models
GPT-LLM BriLLM
Transformer, GPT
Deep Learning SiFu Learning
Machine Learning
@@ 2 SiFu Mechanism
The human brain’s representation and processing of information—particularly language—have been characterized in Huth et al. (2016): semantic features of linguistic elements are distributed across the cerebral cortex, with distinct cortical regions encoding specific semantic content, and this organization exhibits inter-individual consistency. This established neural basis of language contrasts sharply with the representational and processing mechanisms of current GPT-LLMs, which only maintain interpretable semantic mappings for linguistic elements at their input and output interfaces. In contrast, the human brain encodes semantics directly within distributed cortical regions, rendering it a fully interpretable "large model" if conceptualized as a language-processing system.
Furthermore, the brain differs fundamentally from GPT-LLMs and all existing machine learning models in how it regulates input-output dynamics. The latter process a discrete input stream through opaque internal computations to generate a predictive output stream, with input and output restricted to dedicated interfaces. The brain, however, almost certainly operates via a distinct mechanism. While experimental and ethical constraints limit our understanding of its detailed operational processes (i.e., thinking), the distributed cortical encoding of semantics implies that language-based cognition requires all semantic units (e.g., vocabulary) to function as both inputs and outputs. Consequently, the cortical regions subserving these semantic representations must act as bidirectional nodes for information flow, serving as both input and output sites. Conceptualizing the brain as a machine learning model thus necessitates a mechanism enabling all components to operate as simultaneous input and output interfaces. This requirement leads us to hypothesize a dynamic signaling system—likely electrophysiological signals such as EEG—that can activate these distributed semantic regions (or model components) at any time.
Inspired by the supposed neural signaling in the brain, we introduce Signal Fully-connected Flowing (SiFu)—a novel input-output control mechanism over directed graphs that forms the core of BriLLM. This mechanism comprises two core components: (1) a fully connected directed graph with bidirectional node connections; (2) a dynamic weighting system that modulates signal strength based on node correlations. As shown in Figure 2, the SiFu model is a graph with sparsely activated nodes that transmit nominal signals through tensors. Each node (ideally a neuronal layer) represents a concept or word (e.g., nouns, verbs), while edges model inter-node relationships.
@@
where $\text{Id}()$ denotes the token’s output representation (typically a one-hot vector), $e()$ denotes the input representation (commonly called word vectors), and $\theta$ is the set of parameters to be learned by $M$. Notably, such models must integrate prior outputs into the context for prediction, giving rise to ’attention’ mechanisms—common to both RNN and Transformer architectures. Accordingly, the attention-augmented form becomes
@@
In these models, only the input sequence $w_{1},\ldots,w_{i-1}$ and output $w_{i}$ are directly interpretable; the model $M$ and its parameters $\theta$ require dedicated analysis to elucidate their roles in the learning process. Moreover, model size scales with input context length, as $M$ must process the entire sequence $w_{1},\ldots,w_{i-1}$.
To address these limitations, the SiFu learning mechanism introduces a redesigned model and prediction paradigm. SiFu is defined as a directed fully connected graph $G=\{V,E\}$, where $V$ (nodes) and $E$ (edges) form a fully connected structure by default. Each node in $V$ is explicitly mapped to a token in the vocabulary: for a vocabulary $\{w_{1},w_{2},\ldots,w_{n}\}$ (with size $n$), the node set is $V=\{v_{1},v_{2},\ldots,v_{n}\}$. This design inherently renders the model interpretable, as each node’s meaning is user-defined—though node semantics (i.e., associated tokens) are distinct from the graph’s structural properties. We assume nodes and edges have specific, learnable topological structures and parameters $\theta_{V}$ (node’s hyperparameters and parameters) and $\theta_{E}$ (edge’s hyperparameters and parameters) to govern tensor flow $r$, which propagates across edges and through nodes.
Signal transformation at nodes and edges follows: $r^{\prime}=r\oplus v_{i,\theta_{V}}$ after passing through node $v_{i}$, and $r"=r\otimes e_{ij,\theta_{E}}$ after traversing edge $e_{ij}$, where $\oplus$ and $\otimes$ denote predefined tensor operations for node and edge interactions, respectively. For simplicity, autoregressive prediction focuses on correctly forecasting the next token. In SiFu, given an input sequence $w_{1},\ldots,w_{i-1}$ (mapped to nodes $v_{1},\ldots,v_{i-1}$), a signal tensor $r$ initiates at $v_{1}$, propagates sequentially through these input nodes to $v_{i-1}$, and then continues onward. The next predicted node is determined by maximizing signal energy: the correct node $v_{i}$ is the one where $r$ attains the highest energy $\|r\|$ (with $\|\cdot\|$ denoting a custom tensor energy function), such that
@@
However, to predict the token $v_{i}$, relying solely on the information from the signal tensor passed from the last preceding node $v_{i-1}$, while ignoring all previous nodes, can lead to significant prediction bias. Therefore, like traditional generative models, we also use a mechanism of ’attention’ to combine the signal tensors output by each node before the node to be predicted. Based on the energy score of the fused signal tensor, we make the final prediction, and thus the predictive formula that can ’attend’ to the information of all nodes in the previous input sequence is
$v_{i}=\arg\max_{v^{\prime}}\sum_{k=1}^{i-1}\left\|\alpha_{k}*(r\oplus v_{1}\otimes e_{12}\oplus v_{2}\ldots\otimes e_{k-1,k}\oplus v_{k}\otimes e_{k,v^{\prime}}\oplus v^{\prime})\right\|.$
where $\alpha_{k}$ represents a learnable weight for signaling from $v_{k}$ to $v^{\prime}$.
This defines SiFu’s prediction process (also called inference or decoding). During training, model parameters may simply optimized to reward correct node predictions and penalize incorrect ones.
In summary, signal propagation is governed by energy dynamics: signals strengthen (maximizing energy) along ’correct’ paths, and the next node is activated only if it maximizes signal energy, fostering competition between potential paths. In details, given input tokens, a signal initiates at the starting node and propagates sequentially. Outputs correspond to nodes along the path with maximal signal energy relative to alternatives. For example, Figure 2 illustrates "dog $\to$ love $\to$ meat" as the highest-energy path. During training (Figure 2), the ground-truth sequence must yield the highest energy. For "love $\to$ meat," for instance, energy tensors are computed for the target ("meat") and negative samples ("apple," "iron"), with the loss function maximizing energy for "meat" while minimizing it for negatives.
@@
Overall, SiFu’s design as a directed fully connected graph with signal propagation confers two key advantages:
  1. 1.
    Inherent full interpretability: User-defined entities (concepts, tokens, or interpretable units) map directly to specific graph nodes;
  2. 2.
    Unbounded contextual capacity: Prediction is framed as signal propagation through node activations. Because signals propagate freely across nodes, sequence prediction naturally supports arbitrarily long contexts without increasing model size.
@@ 3 BriLLM Formulation
BriLLM implements the SiFu mechanism for language modeling, as shown in Figure 4. Each vocabulary token corresponds to a node, defined by a GeLU-activated neuron layer with a bias $b\in\mathbb{R}^{d_{node}}$, where $d_{node}$ is the node dimension (number of neurons per node). Edges between nodes $u$ and $v$ are modeled as fully connected matrices $W_{u,v}\in\mathbb{R}^{d_{node}\times d_{node}}$, with $W_{u,v}$ and $W_{v,u}$ enabling bidirectional signaling.
Signal propagation begins with an initial tensor:
@@
For a training sequence $u_{1},...,u_{L-1},v_{predict}$, signal $e_{i+1}\in\mathbb{R}^{d_{node}}$ when propagating from $u_{i}$ to $u_{i+1}$ is computed as:
@@
where PE denotes sine-cosine positional encoding. Edge-specific biases $b_{u_{i},u_{i+1}}$ modulate signaling, with $b_{u_{1}}$ (edge-independent) initiating sequences.
To predict the next token, an aggregated signal tensor $\mathcal{E}_{i}\in\mathbb{R}^{d_{node}}$ is computed via a weighted sum of previous signals using learnable weights $\alpha\in\mathbb{R}^{L-1}$:
@@
where $L$ is sequence length, and $\mathcal{A}$ represents softmax-normalized weights. This allows the model to "attend" to all preceding tokens, not just the immediate predecessor. Note here $\mathcal{A}$ represents a rank-1 tensor weight for attending each previous signal. However, we empirically discovered higher ranked weighting tensor may further strengthen the long sequence prediction of BriLLM.
The final signal tensor for predicting token $v$ is:
$E_{u,v}={\rm GeLU}(W_{u_{L-1},v}\mathcal{E}_{L-1}+b_{u_{L-1},v}+PE_{L-1}),$
During inference, the predicted token $v_{predict}$ maximizes the L2 norm of this signal tensor:
$v_{predict}$ $=\mathop{\arg\max}_{v}\|{E_{u,v}}\|_{2}$ (4)
Figure 3: The architecture of BriLLM.
Training BriLLM involves constructing a dedicated neural network for each sequence sample, which is illustrated in Figure 4. The network connects input nodes sequentially, with all potential paths integrated into a final softmax layer that identifies the correct path via a 0-1 ground-truth vector. Cross-entropy loss optimizes this process.
@@ 4 Experiments
We released BriLLM-Chinese and BriLLM-English, with experimental details below.
@@
Datasets
BriLLM-Chinese and BriLLM-English were trained on Chinese and English Wikipedia, respectively (each with over 100M tokens). Long sentences were truncated to 32 tokens, with a 4,000-token vocabulary for both languages.
Implementation Details.
BriLLM is implemented in PyTorch, using sine-cosine positional encoding, GeLU activation, and cross-entropy loss for next-token prediction. Nodes have dimension $d_{node}$ = 32. Training used the AdamW optimizer ($\beta_{1}$= 0.9, $\beta_{2}$ = 0.999, $\epsilon=10^{-8}$) on 8 NVIDIA A800 GPUs for 1.5k steps. The theoretical model size is 512 + 4000 $\times 4000\times$ (32 $\times$ 32 + 32) $\approx$ 16B parameters, with training loss shown in Figure 5.
Sparse Training
BriLLM leverages sparse token co-occurrence: most bigrams are low-frequency or absent, allowing shared parameters for inactive edges. Low-frequency bigrams use a fixed, non-updated matrix, reducing model size to 2B (Chinese) and 1B (English)—13.0% and 5.7% of the original size, respectively (Table 2). This reduces parameters by  90% while accelerating training.
@@
Complexity
Forward computation complexity is O($L\cdot n\cdot d_{node}^{2}$), where $L$ is sequence length, $n$ is vocabulary size, and $d_{node}$ is node dimension.
Case Study
Tables  3 and  4 show decoding results for training and test samples in Chinese and English, respectively.

### v7

@@ Abstract
We present BriLLM, a brain-inspired large language model that fundamentally reimagines machine learning foundations through Signal Fully-connected flowing (SiFu) learning. Addressing core limitations in Transformer-based models—black-box opacity, quadratic complexity, and context-length dependency—BriLLM incorporates two key neurocognitive principles: (1) static semantic mapping where tokens map to specialized nodes analogous to cortical regions, and (2) dynamic signal propagation simulating electrophysiological information flow. This architecture enables three breakthroughs: full model interpretability, context-length independent scaling, and the first global-scale simulation of brain-like processing. Initial 1–2B parameter models demonstrate GPT-1-level generative capabilities with stable perplexity reduction. Scalability analyses confirm feasibility of 100–200B parameter variants processing 40,000-token contexts. BriLLM establishes a new paradigm for biologically grounded AGI development.  [^1] Code at: https://github.com/brillm05/BriLLM0.5  [^2] Models at: https://huggingface.co/BriLLM/BriLLM0.5
@@ 1 Introduction
Artificial general intelligence (AGI) — enabling systems to autonomously acquire and generalize skills across domains—requires a departure from narrow, task-specific paradigms. Prevailing large language models (LLMs), built on Transformer and GPT architectures(Radford et al., 2018), increasingly face recognition of fundamental limitations in achieving AGI (Vaswani et al., 2017). Paradoxically, these frameworks may represent the apex of conventional machine learning (ML) and deep learning (DL) paradigms. Their failure to scale toward AGI thus reflects not architectural flaws, but a critical limitation in the underlying ML/DL foundation: black-box opacity, where only inputs and outputs are interpretable, with internal mechanisms inscrutable.
This core flaw — innate to all traditional ML/DL systems — persists regardless of architectural tweaks, including advancements in attention mechanisms. Even optimal designs cannot address the fundamental opacity of conventional models. It is quite possible that Transformer or GPT has been the best design in terms of the current ML/DL paradigm. Progress toward AGI therefore demands a rewrite of ML’s foundations, creating interpretable, brain-inspired frameworks. This imperative motivates our introduction of Signal Fully-connected flowing (SiFu) learning — a paradigm shift stemming from two insights: recognition that GPT’s AGI bottlenecks are systemic, not architectural; and principles derived from macroscopic brain function Huth et al. (2016). SiFu is not an incremental improvement but a complete reimagining of learning, fundamentally distinct from prior ML/DL work.
Conventional LLMs exacerbate this systemic issue through attention-driven inefficiencies (quadratic complexity with sequence length) and parameter scaling tied to context length — contrasting sharply with the brain, which processes arbitrary context without physical expansion. While prior work has borrowed isolated neural features, none have attempted global simulation of brain-like information processing as a basis for intelligence.
Here, we present three interconnected innovations:
$\bullet$ A novel paradigm replacing traditional ML/DL with brain-inspired approach to semantic representation;
$\bullet$ A generative framework grounded in dynamic energy-maximizing signaling;
$\bullet$ The first global-scale computational model of brain-like semantic and functional mechanisms.
Table 1 situates this work within ML evolution, highlighting SiFu’s divergence from conventional paradigms toward brain-aligned learning.
Table 1: Evolution from machine learning to brain-inspired learning
Level Conventional Brain-inspired
$\uparrow$ Application Task-specific models Generalist AGI systems
Architecture Transformer/GPT BriLLM
Framework Deep learning SiFu learning
Foundation Machine learning Neurocognitive principles
@@ 2 SiFu Mechanism
The human brain offers a blueprint for overcoming conventional ML’s limitations. As Huth et al. (2016) demonstrated, semantic information maps to specific cortical regions consistently across individuals — every brain area contributes to interpretable processing, unlike conventional LLMs where only inputs and outputs are transparent. Additionally, cognition arises from electrophysiological signals (e.g., EEG) propagating across these regions, dynamically activating stored knowledge. These two properties — static distributed semantics and dynamic signaling — define brain function and are absent in current ML/DL and AI.
SiFu (Signal Fully-connected flowing) learning replicates these properties through a directed graph $G=\{V,E\}$, fundamentally redefining learning:
  1. 1.
    Static semantic grounding: Nodes $V$ explicitly map to tokens, mirroring cortical regions encoding specific meanings Huth et al. (2016). This ensures full interpretability across every component, not just inputs/outputs, has semantic meaning.
  2. 2.
    Dynamic signal propagation: Edges $E$ enable bidirectional signaling, modeled after neural electrophysiology. Signals flow along "least resistance" paths, maximizing energy — a proxy for neural pathway strengthening during cognition.
This design addresses conventional ML’s core failures. Prediction relies on signal dynamics, not forward computation in a black box; model size is decoupled from sequence length (like the brain); and every component is interpretable.
@@
where $\text{Id}()$ denotes the token’s output representation (typically a one-hot vector), $e()$ denotes the input representation (commonly called word vectors), and $\theta$ is the set of parameters to be learned by $M$. Here $\text{Id}()$ and $e()$ represent what is only explainable in this DL framework. Notably, such models must integrate prior outputs into the context for prediction, giving rise to ’attention’ mechanisms — common to both RNN and Transformer architectures. Accordingly, the attention-augmented form becomes
@@
In these models, only the input sequence $w_{1},\ldots,w_{i-1}$ and output $w_{i}$ are directly understood; the model $M$ and its parameters $\theta$ require dedicated analysis to elucidate their roles in the learning process. Moreover, model size scales with input context length, as $M$ must process the entire sequence $w_{1},\ldots,w_{i-1}$ once through a unique input port.
To address these limitations, formally, SiFu is defined for a vocabulary $\{w_{1},...,w_{n}\}$ with nodes $V=\{v_{1},...,v_{n}\}$ (each $v_{i}$ mapping to $w_{i}$). Edges $e_{ij}\in E$ between $v_{i}$ and $v_{j}$ govern signal transmission via learnable parameters. A signal tensor $r$ initiates at nodes corresponding to input tokens and propagates through the graph, with transformations at nodes ($\oplus$) and edges ($\otimes$) determined by learnable parameters $\theta_{V}$ (nodes) and $\theta_{E}$ (edges).
In SiFu mechanism, given input tokens $w_{1},...,w_{i-1}$ (mapped to $v_{1},...,v_{i-1}$), the signal $r$ propagates through these nodes. The next token $w_{i}$ is identified as the node $v_{i}$ where the signal attains maximum energy, computed as:
@@
For autoregressive prediction (like GPT), the corresponding maximum energy is instead computed by
$v_{i}=\arg\max_{v^{\prime}}\sum_{k=1}^{i-1}\left\|\alpha_{k}*(r\oplus v_{1}\otimes e_{12}\oplus v_{2}\ldots\otimes e_{k-1,k}\oplus v_{k}\otimes e_{k,v^{\prime}}\oplus v^{\prime})\right\|,$
where $\alpha_{k}$ are learnable weights enabling the model to "attend" to relevant prior nodes, mimicking distributed neural integration.
Figure 2 illustrates SiFu’s operation: during forward propagation (Figure 2), signals flow through nodes, with the next token determined by maximum energy; during training (Figure 2), parameters are optimized to ensure correct paths yield the highest energy, analogous to strengthening neural pathways through learning.
SiFu’s key advantages arise directly from its brain-inspired design:
  1. 1.
    Full interpretability: Every node maps to a token, making semantic processing transparent at all levels — replicating the brain’s distributed interpretability Huth et al. (2016).
  2. 2.
    Unbounded context processing: Like the brain, SiFu processes arbitrarily long sequences without expanding its structure, as signal propagation rather than parameter scaling handles longer inputs.
  3. 3.
    Dynamic signaling: Signal flow mirrors electrophysiological activity, enabling recall and activation patterns analogous to human cognition.
  4. 4.
    Cognitive traceability: Thanks to signal propagation and activation of predictions across interpretable nodes, dynamic prediction behavior is explainable throughout the process, realizing cognitive traceability. Error generation can be localized to specific signal paths (e.g., nodes or edges with abnormal activation), similar to analyzing abnormal brain activity via neuroimaging.
@@ 3 BriLLM Formulation
BriLLM implements the SiFu mechanism to realize a language model that replicates the brain’s macroscopic properties and processing, as shown in Figure 3. Each token corresponds to a node — modeled as a GeLU-activated neuron layer with bias $b\in\mathbb{R}^{d_{node}}$ (where $d_{node}$ is node dimension) — mirroring a cortical region dedicated to specific semantics Huth et al. (2016). Edges between nodes $u$ and $v$ are bidirectional, with matrices $W_{u,v},W_{v,u}\in\mathbb{R}^{d_{node}\times d_{node}}$ enabling signal transmission in both directions, analogous to reciprocal neural connections.
Figure 3: The architecture of BriLLM.
Signal propagation in BriLLM mimics electrophysiological activity, starting with an initial tensor:
@@
For a sequence $u_{1},...,u_{L-1},v_{predict}$, the signal $e_{i+1}\in\mathbb{R}^{d_{node}}$ propagating from $u_{i}$ to $u_{i+1}$ is:
@@
Here, positional encoding (PE) ensures sequence order is preserved, while edge-specific biases modulate signal strength.
To predict the next token, BriLLM integrates signals from all preceding nodes using learnable weights $\alpha\in\mathbb{R}^{L-1}$:
@@
where $\mathcal{A}$ is softmax-normalized to prioritize relevant signals. The final prediction is the node maximizing the energy of the propagated signal:
$v_{predict}$ $=\mathop{\arg\max}_{v}{\rm\|GeLU}(W_{u_{L-1},v}\mathcal{E}_{L-1}+b_{u_{L-1},v}+PE_{L-1})\|_{2}$ (4)
Training BriLLM involves optimizing parameters to maximize signal energy for correct sequences, analogous to strengthening neural pathways through learning. As shown in Figure 4, each training sample constructs a dynamic network reflecting the sequence’s signal flow, with cross-entropy loss rewarding accurate energy-based predictions.
@@ 4 Experiments
BriLLM was designed as a generative large model targeting supervised fine-tuning (SFT) capabilities — distinct from early small pre-trained language models like GPT-1 (which focused on deep representation learning). SiFu’s departure from deep representation learning further distinguishes BriLLM, precluding direct comparisons to GPT-1’s representation learning benchmarks or standard fine-tuning evaluations. Additionally, current computational constraints limit our checkpoints to sub-scale sizes, insufficient for demonstrating GPT-LLM-like emergent abilities (SFT validation is thus future work). Instead, we validate two core generative functions: sequence continuation and stable learning dynamics — sufficient to confirm BriLLM’s design feasibility.
4.1 Setup
Datasets: BriLLM-Chinese and BriLLM-English were trained on Chinese and English Wikipedia (each >100M tokens), with sequences truncated to 32 tokens and a 4,000-token vocabulary. This setup tests the model’s ability to process natural language while maintaining the brain-like property of fixed size regardless of sequence length.
Implementation Details: Implemented in PyTorch, BriLLM uses sine-cosine positional encoding, GeLU activation, and cross-entropy loss. Nodes have dimension $d_{node}$=32 (neurons per node), with edges as 32 $\times$ 32 matrices. Training used the AdamW optimizer ($\beta_{1}$=0.9, $\beta_{2}$=0.999) on 8 NVIDIA A800 GPUs for 1.5k steps. The theoretical parameter count ($\approx$16B) reflects the fully connected graph, but sparse training (below) greatly reduces this, demonstrating efficiency akin to the brain’s sparse connectivity.
Sparse Training: Consistent with the brain’s sparse neural connections, BriLLM leverages low-frequency token co-occurrences to reduce parameters. Low-frequency edges share fixed matrices, reducing size to  2B (Chinese) and  1B (English)—90% smaller than theoretical (Table 2). This mirrors the brain’s ability to reuse neural pathways for infrequent concepts.
4.2 Results
Learning stability: Training loss (Figure 5) shows consistent reduction, confirming effective pattern learning.
Sequence continuation: Tables 3 and 4 demonstrate contextually relevant completions, matching GPT-1’s core generative capability (its most impactful feature, despite original focus on representation learning).
@@ 4.2 Results
4.3 Scalability
BriLLM’s size scales quadratically with node dimension, i.e., O($n^{2}\cdot d_{node}^{2}$), but as a global brain simulation, mature models will not require drastic expansion for diverse AGI tasks (like the human brain). Even with a 40,000-token vocabulary (matching modern LLMs), sparse training limits size to 100–200B parameters — competitive with current models — while maintaining unique context-length independence (O(1) model complexity vs. Transformers’ quadratic O($L^{2}$) scaling, with context length $L$).